OpenStudy (anonymous):

determine the intercepts (x and y) (1-x^2)(x^2-5)

OpenStudy (anonymous):

ok i was giving the function just like this i was wondering since its already factorized if i'm just to solve each individual bracket for x

OpenStudy (anonymous):

yes it is in factored form almost \[1-x^2=(1+x)(1-x)\] so that part has zeros of 1 and-1

OpenStudy (anonymous):

\[x^2-5=(x+\sqrt{5})(x-\sqrt{5})\] which has zeros at \[\pm\sqrt{5}\]

OpenStudy (anonymous):

so you have 4 zeros, \[1,-1,\sqrt{5}, -\sqrt{5}\]

OpenStudy (anonymous):

those are the x - intercepts. so get the y - intercept set x = 0 and get -5 from your eyeballs making the y- intercept (0,-5)

OpenStudy (anonymous):

i dont mutiply because it would still end up being the same thing right?

OpenStudy (anonymous):

oh heck no!

OpenStudy (anonymous):

do not multiply you will only have to factor afterward. to find the zeros you want things in factored form!

OpenStudy (anonymous):

of course you will get the same zeros but you will have to factor to get them so by all means do not multiply out first

OpenStudy (anonymous):

oh ok thanks satlitte greatly appreciated